# iLRM: An Iterative Large 3D Reconstruction Model

### v1

@@ Appendix D Computational Costs of Training
Assuming a hidden dimension of $D=768$, an image resolution of $\times 256$ 256, and a viewpoint resolution of $\times 128$ 128, with a patch size of $\times 8$ 8, the token lengths are computed as $L_{i}=1024$ for the image tokens and $L_{v}=256$ for the viewpoint tokens, based on the experimental configuration used in the RealEstate10K zhou2018re10k dataset.
@@
Appendix E Additional Quantitative Results
Varying baseline range. We compare our model against recent generalizable 3D reconstruction methods chen2025mvsplat ; xu2024depthsplat ; nam2024generative on the RealEstate10K zhou2018re10k dataset, with a particular focus on handling varying degrees of camera overlap ye2024no . These overlap categories are determined using the dense feature matching method RoMA edstedt2024roma . As shown in Tab. 6, our method, which efficiently handles a large number of input viewpoints/images, achieves superior performance compared to existing approaches, especially in challenging cases with small viewpoint overlap (i.e., wide-baseline settings).
@@ Appendix E Additional Quantitative Results
MVSplat chen2025mvsplat 20.37 0.725 0.250 23.81 0.814 0.172 27.47 0.885 0.115 24.01 0.812 0.175
DepthSplat xu2024depthsplat 22.82 0.798 0.193 25.38 0.851 0.145 28.32 0.900 0.104 25.59 0.852 0.145
Gen-Den nam2024generative 21.10 0.744 0.234 24.57 0.828 0.162 28.26 0.895 0.108 24.77 0.826 0.164
Ours $(2,F,F)$ 23.66 0.809 0.189 26.20 0.858 0.145 28.82 0.901 0.110 26.33 0.858 0.146
Ours $(4,H,F)$ 27.45 0.883 0.130 28.83 0.903 0.110 30.30 0.921 0.094 28.91 0.903 0.111
Ours-MC $(4,H,F)$ 27.22 0.878 0.134 28.57 0.899 0.114 30.02 0.918 0.098 28.66 0.899 0.114
Ours $(8,H,F)$ 28.87 0.904 0.113 29.96 0.918 0.099 31.20 0.932 0.085 30.05 0.919 0.098
Ours-MC $(8,H,F)$ 28.72 0.903 0.114 29.73 0.916 0.101 30.85 0.929 0.088 29.80 0.916 0.101
Table 6: Quantitative comparisons on the RE10K zhou2018re10k dataset under varying view overlap conditions.
Same number of Gaussians. We also validate the strength of our decoupling strategy in leveraging high-resolution images as visual cues while generating efficient and compact low-resolution 3D Gaussians. As discussed in our motivation, previous methods charatan2024pixelsplat ; chen2025mvsplat ; zhang2025gs-lrm ; xu2024depthsplat ; nam2024generative require downsampling the input images to reduce the number of generated Gaussians, inherently coupling image resolution with representation density. To demonstrate the flexibility of our approach, we conduct an experiment in which all methods generate the same number of Gaussians using 4 viewpoints at half resolution. Specifically, the baseline methods chen2025mvsplat ; xu2024depthsplat follow a $(4,H,H)$ configuration, where both the number of viewpoints and image resolution are reduced. In contrast, our method adopts a $(4,H,F)$ setting, where we preserve high-resolution image inputs while generating low-resolution Gaussians, thanks to our decoupled design. As shown in Tab. 7, our method surpasses the baselines in performance while requiring fewer computational resources in training, and faster inference speed, highlighting the practical advantages of our design. This result demonstrates the efficiency and the representational ability of our architecture, which effectively utilizes high-resolution visual cues, leading to superior reconstruction quality under the same output density without requiring expensive hardware. To train the baseline methods effectively with a large batch size (similar to ours), we run them on a single H100 GPU. Our method and MVSplat chen2025mvsplat are trained with a batch size of 16, while DepthSplat xu2024depthsplat is trained with a batch size of 12 due to memory constraints.
@@
MVSplat chen2025mvsplat 12 H100 (1) 27.53 0.889 0.116 65,536 0.048 0.65
DepthSplat xu2024depthsplat 354 H100 (1) 28.08 0.898 0.107 65,536 0.062 2.49
Ours 185 RTX 4090 (1) 28.96 0.902 0.111 65,536 0.027 1.22
Ours 185 RTX 4090 (2) 29.50 0.911 0.104 65,536 0.027 1.22
Ours 185 RTX 4090 (4) 29.98 0.918 0.098 65,536 0.027 1.22
Table 7: Quantitative comparisons under the same number of Gaussians on the RE10K  zhou2018re10k dataset. Inference time and memory consumption are measured only during the Gaussian generation stage, excluding the rendering process on a single RTX 4090 GPU.
@@
Ours $(8,H,F)$ 30.15 0.920 0.097 131,072
Ours $(4,H,F)$ 28.96 0.902 0.111 65,536
Ours $(8,Q,F)$ 27.21 0.864 0.154 32,768
Ours $(4,Q,F)$ 26.25 0.839 0.180 16,384
Table 8: Quantitative comparisons of different viewpoint configurations on the RE10K zhou2018re10k dataset. $Q$ denotes quarter resolution compared to the original image resolution.
Inference time and memory usage. We benchmark the inference time and peak GPU memory usage for generating Gaussians with varying numbers of input viewpoints/images. We use an image resolution of 256 $\times$ 448 and a viewpoint resolution of 128 $\times$ 224, following the DL3DV dataset experiment settings, and conduct measurements on a single RTX 4090 GPU. Note that, 50 target images are allocated as a bias in accordance with the evaluation setting presented in the main paper.
# of frames 2 4 6 8 12 16 32 64 128 256
Time (S) 0.027 0.028 0.031 0.039 0.054 0.081 0.181 0.458 1.285 4.065
Memory (GB) 1.29 1.36 1.40 1.45 1.62 1.76 2.31 3.42 5.65 10.11
Table 9: Inference time and memory usage with varying numbers of input viewpoints/images.
Appendix F Additional Qualitative Results
We present additional qualitative results in Fig. 11 and Fig. 12 corresponding to the RealEstate10K (RE10K) zhou2018re10k and DL3DV ling2024dl3dv datasets, respectively. Also, Fig. 13 presents high-resolution visualizations from the DL3DV dataset.
Figure 11: Qualitative comparison of novel view synthesis on the RE10K dataset.
Figure 12: Qualitative comparison of novel view synthesis on the DL3DV dataset.
Figure 13: High-resolution novel view synthesis on the DL3DV dataset.

### v2

@@ Appendix F Computational Costs of Training
Assuming a hidden dimension of $D=768$, an image resolution of $256\times 256$, and a viewpoint resolution of $128\times 128$, with a patch size of $8\times 8$, the token lengths are computed as $L_{i}=1024$ for the image tokens and $L_{v}=256$ for the viewpoint tokens, based on the experimental configuration used in the RealEstate10K Zhou et al. (2018) dataset.
@@
Appendix G Additional Quantitative Results
Varying baseline range. We compare our model against recent generalizable 3D reconstruction methods (Chen et al., 2025; Xu et al., 2025; Nam et al., 2025) on the RealEstate10K (Zhou et al., 2018) dataset, with a particular focus on handling varying degrees of camera overlap (Ye et al., 2024). These overlap categories are determined using the dense feature matching method RoMA (Edstedt et al., 2024). As shown in Tab. 7, our method, which efficiently handles a large number of input viewpoints/images, achieves superior performance compared to existing approaches, especially in challenging cases with small viewpoint overlap (i.e., wide-baseline settings).
@@ Appendix G Additional Quantitative Results
MVSplat (Chen et al., 2025) 20.37 0.725 0.250 23.81 0.814 0.172 27.47 0.885 0.115 24.01 0.812 0.175
DepthSplat (Xu et al., 2025) 22.82 0.798 0.193 25.38 0.851 0.145 28.32 0.900 0.104 25.59 0.852 0.145
Gen-Den (Nam et al., 2025) 21.10 0.744 0.234 24.57 0.828 0.162 28.26 0.895 0.108 24.77 0.826 0.164
Ours $(2,F,F)$ 23.82 0.813 0.184 26.54 0.864 0.139 29.43 0.910 0.103 26.70 0.864 0.140
Ours $(4,H,F)$ 27.65 0.887 0.127 29.13 0.908 0.108 30.73 0.926 0.092 29.22 0.908 0.108
Ours-MC $(4,H,F)$ 27.41 0.882 0.131 28.87 0.904 0.111 30.44 0.927 0.095 28.96 0.904 0.112
Ours $(8,H,F)$ 29.44 0.912 0.106 30.51 0.925 0.093 31.77 0.937 0.080 30.61 0.925 0.092
Ours-MC $(8,H,F)$ 28.92 0.906 0.112 29.97 0.920 0.098 31.15 0.932 0.086 30.05 0.920 0.098
Table 7: Quantitative comparisons on the RE10K dataset under varying view overlap conditions.
Same number of Gaussians. We also validate the strength of our decoupling strategy in leveraging high-resolution images as visual cues while generating efficient and compact low-resolution 3D Gaussians. As discussed in our motivation, previous methods (Charatan et al., 2024; Chen et al., 2025; Zhang et al., 2025; Xu et al., 2025; Nam et al., 2025) require downsampling the input images to reduce the number of generated Gaussians, inherently coupling image resolution with representation density. To demonstrate the flexibility of our approach, we conduct an experiment in which all methods generate the same number of Gaussians using 4 viewpoints at half resolution. Specifically, the baseline methods (Chen et al., 2025; Xu et al., 2025) follow a $(4,H,H)$ configuration, where both the number of viewpoints and image resolution are reduced. In contrast, our method adopts a $(4,H,F)$ setting, where we preserve high-resolution image inputs while generating low-resolution Gaussians, thanks to our decoupled design. As shown in Tab. 8, our method surpasses the baselines in performance while requiring fewer computational resources in training, and faster inference speed, highlighting the practical advantages of our design. This result demonstrates the efficiency and the representational ability of our architecture, which effectively utilizes high-resolution visual cues, leading to superior reconstruction quality under the same output density without requiring expensive hardware. To train the baseline methods effectively with a large batch size (similar to ours), we run them on a single H100 GPU. Our method and MVSplat (Chen et al., 2025) are trained with a batch size of 16, while DepthSplat (Xu et al., 2025) is trained with a batch size of 12 due to memory constraints.
@@
MVSplat (Chen et al., 2025) 12 H100 (1) 27.53 0.889 0.116 65,536 0.048 0.65
DepthSplat (Xu et al., 2025) 354 H100 (1) 28.08 0.898 0.107 65,536 0.062 2.49
Ours 185 RTX 4090 (1) 29.24 0.907 0.109 65,536 0.027 1.22
Ours 185 RTX 4090 (2) 29.82 0.916 0.101 65,536 0.027 1.22
Table 8: Quantitative comparisons under the same number of Gaussians on the RE10K dataset. Inference time and memory consumption are measured only during the Gaussian generation stage, excluding the rendering process on a single RTX 4090 GPU.
@@
Ours $(8,H,F)$ 30.39 0.923 0.095 131,072
Ours $(4,H,F)$ 29.24 0.907 0.109 65,536
Ours $(8,Q,F)$ 27.36 0.868 0.152 32,768
Ours $(4,Q,F)$ 26.40 0.843 0.177 16,384
Table 9: Quantitative comparisons of different viewpoint configurations on the RE10K dataset. $Q$ denotes quarter resolution compared to the original image resolution.
Appendix H Additional Qualitative Results
We present additional qualitative results in Fig. 12 and Fig. 13 corresponding to the RealEstate10K (RE10K) (Zhou et al., 2018) and DL3DV (Ling et al., 2024) datasets, respectively. Also, Fig. 14 presents high-resolution visualizations from the DL3DV dataset.
Figure 12: Qualitative comparison of novel view synthesis on the RE10K dataset ($\times256$).
[…]